OpenStudy (anonymous):

I want to construct a rectangular storage container that has a square base, a closed top, and a volume of 9ft^3. If the material for the base costs $5/ft^2 and the material for the sides and to costs $4/ft^2, what is the cost of the cheapest such container?

OpenStudy (anonymous):

What i've got so far is are of base is w*l, cost of was is 5wl cost of front is 4lh cost of back is 4lh cost of left side is 4wh cost o right side is 4wh .... total cost is 5wl+2(4lh)+2(4wh) =5wl+8lh+8wl

OpenStudy (anonymous):

4wl+8lh+8wh (typo)

OpenStudy (campbell_st):

so you are combining volume and surface area if the base has dimensions of l x l the area is l^2 and let the height = h the volume is \[V = l^2 \times h\] and you know the volume is 9 so \[9 = l^2 \times h\] this formula can be changed to make h the subject \[h = \frac{9}{l^2}\] the surface areas top and bottom area are the same \[A =2 l^2\] the 4 sides have length l and width h so there areas are \[A = 4 \times l \times h\] So the cost of the box is \[C = 4(2l^2) + 5(4l \times \frac{9}{l^2})\] so the cost can be simplified to \[C(l) = 8l^2 + \frac{180}{l}\] I'd then differentiate with respect to l \[C'(l) = 16l - \frac{180}{l^2}\] set C'(l) = 0 and some for l. this will give the value of l that gives the minimum cost. you would then need to substitute to find the height value. Hope this makes sense.... and you have done calculus.

OpenStudy (anonymous):

Thank you... what I don't understand is how you put it all into a formula. I understand how to label it, I just want to understand how you went about making the forumla

OpenStudy (campbell_st):

well you have the right idea with your work so far, its just the base, and lid, are squares so you are only dealing with 2 variables, length and height. so you are using the volume information to express combined surface area with 1 variable, length. For most of these types of questions, it just practice and carefully reading... and the key yo 1 formula is you are asked for minimum cost... so that has to be the equation, or formula.

OpenStudy (anonymous):

I'm sorry, but I still do not quite understand.

OpenStudy (anonymous):

Can you explain it in terms of v=l*w*h with the information that I have?

OpenStudy (anonymous):

Like the formula I made is 5wl+8lh+8wh .... what i'm asking is how could I find the length, width and heigh using the v=lwh formula?

OpenStudy (anonymous):

(i'm trying to find the domain and critical numbers)

OpenStudy (campbell_st):

ok... so the square base  you have 2 squares, top and base of the box. the sides  you have 4 of these recatangles so the total area of the surfaces are \[A = 2l^2 + 4 \times l \times h\] now to make times simpler... you know some volume information with a square base the volume of a prism is V = Base area x height so \[9 = l^2 \times h\] making h the subject you get \[h = \frac{9}{l^2}\] and lastly changing the Total surface area to a cost equation $4 per ft^2 for base and top $5 for sides \[C = 4(2l^2) + 5(4lh)\] use the value for h above and substitute it into the cost equation.